List of the Shorthands for Journal Names in Various .mbs Files

Patrick W. Daly

This is part of the makebst or custom-bib bundle

The shorthands here may be used in BibTEX database (.bib) files in place of the full journal name. For example, instead of

journal = "Phys. Rev.",

it is sufficient to give

journal = pr,

The advantage of this is not only the convenience of typing a shorter form, but also the automatic switching between the abbreviated (here “Phys. Rev.”) and full forms (“Physical Review”) depending on the bibliographic style file. Thus the one database entry is reformatted automatically.

There is a set of standard shorthands for computing science included in the

original plain.bst, abbrv.bst, unsrt.bst, and alpha.bst files. These are

built into merlin.mbs and are always present in every .bst file generated from it.

Whether or not the shorthand is present in your .bst file depends on the inclusion of the proper .mbs source file during its creation with the makebst program. Check the heading to the .bst file to find out what source files went into its generation.

You may make up your own list of journal names, abbreviated and full, by using one of the existing .mbs files as a model. Include its name in shorthnd.ins to obtain a listing of its shorthands too.

File: merlin.mbs, version: (all)

acmcs  ACM Computing Surveys
acta  Acta Informatica
cacm  Communications of the ACM

ibmjrd IBM Journal of Research and

Development

ibmsj  IBM Systems Journal

ieeese IEEE Transactions on Software

Engineering

ieeetc  IEEE Transactions on Computers

ieeetcad IEEE Transactions on Computer-Aided Design of Integrated Circuits

ipl  Information Processing Letters
jacm  Journal of the ACM

jcss Journal of Computer and System

Sciences

scp  Science of Computer Programming
sicomp  SIAM Journal on Computing

tocs ACM Transactions on Computer

Systems

tods ACM Transactions on Database

Systems

tog  ACM Transactions on Graphics

toms ACM Transactions on Mathematical

Software

toois ACM Transactions on Office

Information Systems

toplas ACM Transactions on Programming

Languages and Systems

tcs  Theoretical Computer Science

File: physjour.mbs, version: 1997/09/15 2.0b (PWD)

aa  Astronomy and Astrophysics
aph  Acta Physica
advp  Advances in Physics
ajp  American Journal of Physics
ajm  American Journal of Mathematics
amsci  American Scientist
anofd  Annals of Fluid Dynamics
am  Annals of Mathematics
ap  Annals of Physics

adp  Annalen der Physik
ao  Applied Optics
apl  Applied Physics Letters
app  Astroparticle Physics
apj  Astrophysical Journal

baps Bulletin of the American Physical

Society

cmp Communications in Mathematical

Physics

cpam Communications on Pure and Applied

Mathematics

cppcf Comments on Plasma Physics and

Controlled Fusion

cpc  Computer Physics Communications
cqg  Classical and Quantum Gravity

cra Comptes Rendus des Séances de

l’Académie des Sciences A

fed  Fusion Engineering and Design
ft  Fusion Technology
grg  General Relativity and Gravitation
ieeens  IEEE Transactions on Nuclear Science
ieeeps  IEEE Transactions on Plasma Science

ijimw International Journal of Infrared and

Millimeter Waves

ip  Infrared Physics
irp  Infrared Physics
jap  Journal of Applied Physics

jasa Journal of the Acoustical Society of

America

jcp  Journal of Computational Physics
jetp  Soviet Physics–JETP
jfe  Journal of Fusion Energy
jfm  Journal of Fluid Mechanics
jmp  Journal of Mathematical Physics
jne  Journal of Nuclear Energy

jnec Journal of Nuclear Energy, Part C: Plasma Physics, Accelerators, Thermonuclear Research

jnm  Journal of Nuclear Materials
jpc  Journal of Physical Chemistry
jpp  Journal of Plasma Physics

jpsj Journal of the Physical Society of

Japan

jsi  Journal of Scientific Instruments

jvst Journal of Vacuum Science and Technology

nat  Nature

nature  Nature

nedf Nuclear Engineering and

Design/Fusion

nf  Nuclear Fusion
nim  Nuclear Instruments and Methods

nimpr Nuclear Instruments and Methods in

Physics Research

np  Nuclear Physics
npb  Nuclear Physics B
nt/f  Nuclear Technology/Fusion
inc  Il Nuovo Cimento
nc  Nuovo Cimento
pf  Physics of Fluids
pfa  Physics of Fluids A: Fluid Dynamics
pfb  Physics of Fluids B: Plasma Physics
pl  Physics Letters
pla  Physics Letters A
plb  Physics Letters B
prep  Physics Reports

pnas Proceedings of the National Academy

of Sciences of the USA

pp  Plasma Physics
ppcf  Plasma Physics and Controlled Fusion

phitrsl Philosophical Transactions of the

Royal Society of London

prl  Physical Review Letters
pr  Physical Review
physrev Physical Review
pra  Physical Review A
prb  Physical Review B
prc  Physical Review C
prd  Physical Review D
pre  Physical Review E
ps  Physica Scripta

procrsl Proceedings of the Royal Society of

London

rmp  Reviews of Modern Physics
rsi  Review of Scientific Instruments

science Science

sciam  Scientific American
sam  Studies in Applied Mathematics
sjpp  Soviet Journal of Plasma Physics
spd  Soviet Physics–Doklady
sptp  Soviet Physics–Technical Physics
spu  Soviet Physics–Uspeki

File: geojour.mbs, version: 1997/09/15 2.0b (PWD)

aisr  Advances in Space Research
ag  Annales Geophysicæ
anigeo  Annali di Geofisica
angl  Annals of Glaciology
andmet  Annalen der Meteorologie
andgeo  Annales de Geophysique
andphy  Annales de Physique

afmgb Archiv für Meteorologie, Geophysik

und Bioklimatologie

atph  Atmósphera
aao  Atmosphere-Ocean
atenv  Atmospheric Environment

aujag Australian Journal of Agricultural

Research

aumet  Australian Meteorological Magazine
blmet  Boundary-Layer Meteorology

bams Bulletin of the American

Meteorological Society

cch  Climatic Change
cdyn  Climate Dynamics
cbul  Climatological Bulletin
cap  Contributions to Atmospheric Physics
dsr  Deep-Sea Research
dhz  Deutsche Hydrographische Zeitschrift
dao  Dynamics of Atmospheres and Oceans
eco  Ecology
envres  Environmental Research
envst  Environmental Science and Technology
ecms  Estuarine and Coastal Marine Science
geoint  Geofisica Internationale
geopub  Geofysiske Publikasjoner
geogeo  Geologiya i Geofizika

gafd Geophysical and Astrophysical Fluid

Dynamics

gfd  Geophysical Fluid Dynamics
geomag  Geophysical Magazine
georl  Geophysical Research Letters
grl  Geophysical Research Letters
ga  Geophysica
gs  Geophysics

ieeetap IEEE Transactions on Antennas and Propagation ijawp International Journal of Air and Water

Pollution

ijc  International Journal of Climatology

ijrs International Journal of Remote

Sensing

jam  Journal of Applied Meteorology

jaot Journal of Atmospheric and Oceanic

Technology

jatp Journal of Atmospheric and Terrestrial

Physics

jce  Journal of Climate

jcam Journal of Climate and Applied

Meteorology

jcm  Journal of Climate Meteorology
jcy  journal of Climatology
jgr  Journal of Geophysical Research
jga  Journal of Glaciology
jh  Journal of Hydrology
jmr  Journal of Marine Research

jmrj Journal of Meteorological Research,

Japan

jm  Journal of Meteorology
jpo  Journal of Physical Oceanography
jra  Journal de Recherches Atmospheriques
jaes  Journal of the Aeronautical Sciences

japca Journal of the Air Pollution Control

Association

jas  Journal of the Atmospheric Sciences

jmts Journal of the Marine Technology

Society

jmsj Journal of Meteorological Society of

Japan

josj Journal of the Oceanographical

Society of Japan

jwm  Journal of Weather Modification
lao  Limnology and Oceanography
mwl  Mariners Weather Log
mau  Mausam
meteor  “Meteor” Forschungsergebnisse
map  Meteorology and Atmospheric Physics
metmag  Meteorological Magazine
metmon  Meteorological Monographs
metrun  Meteorologische Rundschau
metzeit Meteorologische Zeitschrift
metgid  Meteorologiya i Gidrologiya
mwr  Monthly Weather Review

nwd  National Weather Digest

nzjmfr New Zealand Journal of Marine and

Freshwater Research

npg  Nonlinear Processes in Geophysics
om  Oceanography and Meteorology
ocac  Oceanologica Acta

oceanus Oceanus

paleoc  Paleoceanography
pmg  Papers in Meteorology and Geophysics

ppom Papers in Physical Oceanography and

Meteorology

physzeit Physikalische Zeitschrift
pps  Planetary and Space Science
pss  Planetary and Space Science
pag  Pure and Applied Geophysics

qjrms Quarterly Journal of the Royal

Meteorological Society

quatres Quaternary Research
rsci  Radio Science
rse  Remote Sensing of the Environment
rgeo  Reviews of Geophysics

rgsp Reviews of Geophysics and Space

Physics

rdgeo  Revista de Geofisica
revmeta Revista Meteorologica
ssr  Space Science Review

tellus  Tellus

tac  Theoretical and Applied Climatology

tagu Transactions, American Geophysical

Union (EOS)

wrr  Water Resources Research

weather Weather

wafc  Weather and Forecasting
ww  Weatherwise

wmob World Meteorological Organization

Bulletin

zeitmet Zeitschrift für Meteorologie

File: photjour.mbs, version: 1997/09/15 2.0a (PWD)

appopt  Applied Optics
bell  Bell Systems Technical Journal
ell  Electronics Letters
jasp  Journal of Applied Spectroscopy
jqe  IEEE Journal of Quantum Electronics

jlwt  Journal of Lightwave Technology
jmo  Journal of Modern Optics

josa Journal of the Optical Society of

America

josaa Journal of the Optical Society of

America A

josab Journal of the Optical Society of

America B

jdp  Journal de Physique
oc  Optics Communications
ol  Optics Letters
phtl  IEEE Photonics Technology Letters

pspie Proceedings of the Society for

Photo-Optical Instruments Engineers

sse  Solid-State Electronics
sjot  Soviet Journal of Optical Technology
sjqe  Soviet Journal of Quantum Electronics

sleb Soviet Physics–Lebedev Institute

Reports

stph  Soviet Physics–Technical Physics
stphl  Soviet Technical Physics Letters
vr  Vision Research
zph  Zeitschrift für Physik
zphb  Zeitschrift für Physik B
zphd  Zeitschrift für Physik D
CLEO  Conf. on Lasers and Electro-Optics
ASSL  Advanced Solid State Lasers
OSA  Optical Society of America